OpenStudy (anonymous):

For an amortized loan, you are given the loan, annual interest rate, and length of the loan. Find the monthly payment necessary to pay off the loan. Do not round until the final answer. Then, round to the nearest cent. amount: $9,000; rate: 9.6%; time: 3 years

OpenStudy (anonymous):

do you know the formula for this?

OpenStudy (anonymous):

i think i remember it try \[\frac{9000\times \frac{.096}{12}}{1-\left(1+\frac{.096}{12}\right)^{-36}}\]